please help me solve this equation.c+4=-14.it says check your solution
Found 2 solutions by Theo, unlockmath:
Answer by Theo(13342)

c + 4 = -14
subtract 4 from both sides of the equation to get:
c = -18

to check your equation, replace c with -18 in your equation to see if it is true.

your original equation is c + 4 = -14

replace c with -18 to get:

-18 + 4 = -14

add 18 to both sides of this equation to get:

4 = -14 + 18

simplify and combine like terms to get:

4 = 4

since this is true (4 really does equal to 4), then your value of c = -18 is good.


Answer by unlockmath(1688)

Hello,
Here's what you want to do with equations. Get the letter (in this case c all alone on one side of the equal sign. To do that, you subtract 4 on both sides of the equation.
C + 4 - 4 = -14 - 4
You see what happens? +4 minus 4 is 0 so now you have C all by itself.
So here's what the answer is:
C = - 18
If you are confused how I got -18. Think of it like this. If the temperature is
14 below zero and it got 4 degrees colder, what would be the temperature?
That's right -18
Now plug the -18 into the original equation:
-18 + 4 = -14
Then doing the rest of the calculation you see

-14 = -14
RJ Toftness
Author of "Unlock the Mystery to Math"
www.math-unlock.com